OpenStudy (anonymous):

which set of numbers could be the lengths of the sides of a right triangle ? 10,24,26 12,16,30 3,4,6 4,7,8

OpenStudy (anonymous):

10,24,26

OpenStudy (anonymous):

how did you get that ?

OpenStudy (anonymous):

can you explain it to me ?

OpenStudy (amistre64):

in order to create a triangle, you have to have sides long enough to reach each other

OpenStudy (anonymous):

yess continue . I am reading

OpenStudy (amistre64):

if we put the longest side for our base; and attach the shorter sides to each end; then the short sides have to at least add up to the longer side

OpenStudy (amistre64):



OpenStudy (anonymous):

yes that makes sense

OpenStudy (anonymous):

okay so in this problem how do I figure it out add all the short sides in the set of numbers ? but none of them equal the longest side ?

OpenStudy (amistre64):



OpenStudy (amistre64):

IF short + short is greater than long; you can make a triangle

OpenStudy (amistre64):

add up the smaller values and see if they are bigger than the large value

OpenStudy (amistre64):

the only one the does NOT make a triangle is: 12, 16, 30 12+16 = 28 is less than 30

OpenStudy (anonymous):

okay let me try

OpenStudy (anonymous):

yes that is right but the other ones are triangles so how do we know that number 1 is a right triangle ?

OpenStudy (amistre64):

.... i overlooked the "right" part :)

OpenStudy (anonymous):

not to butt in (ok to butt in) the question asks about right triangles

OpenStudy (anonymous):

Sorry I should've explained: Since it's a right triangle you can check for pythagorean triplets

OpenStudy (anonymous):

your job is to check that \(a^2+b^2=c^2\) if so, the answer is yes if not, the answer is no

OpenStudy (anonymous):

10,24,26 satisfies that

OpenStudy (amistre64):

short^2 + short^2 = long^2 is another way to see it

OpenStudy (anonymous):

thank you all so much for your help this was greattt! now I understand how to do this problem !